Staff:

LBTO

Arcetri

Summary:

System check with teh existing pupils found last day.

Close loop peformances wityh disturbance.

interdiatnce measurements.

Details:

All hand meeting.

luci troubleshooting.

late start.

power on:
  • OK
12:00 am

Test preset OK.

Luci calibraiton unit in the beam. Request to remove it.

Luci-cal_unit_beam.jpg

No light on the CCD47.

1:40 PM There was an error and James R worked on it. Sent email to olga and barry to see if the Luci Cal unit can be parked.

1:45 PM luci Cal unit park

Restart the preset.

PUPIL: ID: 76 FILE:20211110_215148

BIN2

TN:20211112_210914

BIN4

PUPIL:20200916_173858

split x=0, y=1

X=0

Y=3

SPLIT X=0,Y=2

PUPIL=20211112_212830

ROTATION test

START 2:41PM

0->360

2:52PM

First pupil approach for Bin4.
  • shift x=-1 and y=2

Handover

stop 3:30 PM

-- JuanCarlosGuerra - 12 Nov 2021
Topic revision: r1 - 12 Nov 2021, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback